# DFT-s-OFDM-based On-Off Keying for Low-Power Wake-Up Signal

### v1

@@ V-A1 Frequency-Repetition
Starting from the rectangular-like OOK design discussed so far, a generic method to achieve higher frequency diversity is to repeat its subcarrier coefficients on a wider bandwidth. This can be achieved in (7) by selecting $N_{\rm pulse}\ll N_{\rm sc}$, i.e. modulating less, and more spread-apart, DFT-s-OFDM pulses, leading inevitably to a comb-like signal shape. In order to guarantee a good ON/OFF energy split, we first apply (18) to the intermediate narrow band of $N_{\rm pulse}$ subcarriers, i.e. setting $\Phi=\frac{\pi(2L+N_{\rm pulse}-1)}{N_{\rm pulse}}$, before repeating in (7) to populate the $N_{\rm sc}$ subcarriers. An example of the obtained waveforms and power distribution over subcarrier indices is shown on Fig. 10 with $N_{\rm sc}=144$ and $N_{\rm bit}=4$. No FDSS is used here other than a linear phase for time-shifting the signal by half a pulse. As it can be seen the ON symbols cannot be flattened since the main lobes of neighboring pulses simply do not overlap, but by selecting a proper $(\Phi,L)$, the OFF symbols can be guaranteed to be close to zero. Similarly to Fig. 6, variations over $(\Phi,L)$ can, in fact, be tolerated with little effect on the signal shape until high envelope lobes are created in the OFF symbols, similar to the green curve of Fig. 8. The drawbacks of such signals are that it may be less robust to lower sampling rate if considered, and may also increase the PAPR of the overall OFDM transmission (depending on concurrent data constellation).
@@ V-A2 Phase Scrambling in Overlaid Sequence
Methods for varying the phase of the overlaid sequence $r_{0}[k]$ in order to flatten the spectrum were considered in [23]. While by scrambling the phase of the spreading sequence, the envelope shape of the ON symbol may become less controlled, the ON/OFF energy split can still be expected to remain good as it is unlikely that many distant sidelobes in the OFF symbols would coherently combine (cf. $O_{p}[n]$ in (12) for samples outside of $[pN_{\rm seg},\,(p+1)N_{\rm seg}-1]$).
As shown in Corr. 1, the FD power distribution depends on two sequences: $R_{0}[k]$ from the overlaid sequence, and $B[k]$ from the bits which inherently has a repetition structure such that its power is already well-distributed across the band. We can formalize the following.
@@
With Manchester coding [^5] Without Manchester coding, $E[|B[k]|^{2}]=\frac{N_{\rm bit}(N_{\rm bit}+1)}{4}$ for $k=(0\,{\rm mod}\,N_{\rm bit})$ and $E[|B[k]|^{2}]=\frac{N_{\rm bit}}{4}$ otherwise; and more than half, precisely $\frac{(N_{\rm bit}+1)}{N_{\rm bit}}$, of the total average power is allocated on the comb., half of the total power of $\{D_{0}[k]\}_{k=0}^{N_{\rm pulse}-1}$ is always allocated on a comb $\{k=k^{\prime}N_{\rm bit}\}_{k^{\prime}=0}^{N_{\rm seg}-1}$. On average, the power distribution can be verified to be $E[|D_{0}[k]|^{2}]=|R_{0}[k]|^{2}E[|B[k]|^{2}]$ with
$E[|B[k]|^{2}]=\begin{cases}\frac{N_{\rm bit}^{2}}{4}\quad\quad{\rm for}\;\;k=(0\,{\rm mod}\,N_{\rm bit})\\ \frac{N_{\rm bit}}{4}\left(1-\cos\frac{2\pi k}{N_{\rm bit}}\right)\quad{\rm otherwise}\end{cases}.$ (27)
While having half of the power distributed on this comb is independent of $R_{0}[k]$, the power level fluctuation on it depends on $R_{0}[k]$. More precisely, $R_{0}[k]$ is an interpolation of the DFT of $r_{0}[l]$, $R_{0}^{\prime}[k]=\sum_{l=0}^{N_{\rm seg}-1}r_{0}[l]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}kl}$, as
$\!\!\!R_{0}[k]=\begin{cases}R^{\prime}_{0}[k/N_{\rm bit}]\quad\quad\quad{\rm for}\;\;(k\,{\rm mod}\,N_{\rm bit})=0\\ \displaystyle\frac{1}{N_{\rm seg}}\!\!\!\sum_{h=0}^{N_{\rm seg}-1}\!\!\!R^{\prime}_{0}[h]I[k-N_{\rm bit}h]\quad{\rm otherwise}\!\!\!\end{cases}$ (28)
where $I[k]=e^{-{\rm j}\frac{\pi(N_{\rm seg}-1)k}{N_{\rm pulse}}}\frac{\sin\left(\frac{\pi k}{N_{\rm bit}}\right)}{\sin\left(\frac{\pi k}{N_{\rm pulse}}\right)}$, and the power level on the comb $\{k=k^{\prime}N_{\rm bit}\}_{k^{\prime}=0}^{N_{\rm seg}-1}$ is completely controlled by the DFT of $r_{0}[l]$.
@@
Randomizing the phase of $r_{0}[m]$ as considered in [23] would equalize on average the power of its DFT coefficients as $E[|R_{0}[k]|^{2}]=N_{\rm seg}$, which combined with Lem. 4 gives
@@
With random overlaid sequence such that $E[r_{0}[m]r_{0}^{*}[m^{\prime}]]=\delta_{m,m^{\prime}}$, half of the total average power of $\{D_{0}[k]\}_{k=0}^{N_{\rm pulse}-1}$ is equally distributed on a $N_{\rm bit}$-spaced comb.
Therefore, this method could ensure a well-spread energy along the band. The drawbacks are that the instantaneous signal is not controlled, and therefore, its knowledge could not be exploited by receivers with higher-end capabilities.
@@
The approach retained by 3GPP is to use a ZC sequence [23] which is a constant-envelope zero autocorrelation (CAZAC) sequence. A well-known property of a CAZAC sequence is that its DFT is also of constant magnitude [24], such that $|R_{0}^{\prime}[h]|^{2}=N_{\rm seg}$ for all $h$ in (28). Combined with Lem. 4 this gives the same result as in Corr. 4 but for each OFDM symbol without averaging.
@@
With a CAZAC overlaid sequence, half of the total power of $\{D_{0}[k]\}_{k=0}^{N_{\rm pulse}-1}$ is equally distributed on a regular $N_{\rm bit}$-spaced comb.
@@
Comparing to (18) for flattening ON symbols, the closer $\Delta_{\theta}[m]$ would be constant and close to $\Phi\approx\pi$ in this case, the better the DFT-s-OFDM pulses are expected to constructively combine. The deviations of this linear phase increment $\Delta_{\theta}$ from a constant increment of $\pi$ are shown on Fig. 11(e). With roots closer to $(0\mod N_{\rm seg})$, e.g. $u=1$ or $u=(N_{\rm seg}-1)$, the increment changes slower than with ‘middle roots’, e.g. $u=(N_{\rm seg}/2+1)$, which is better for pulse combining. The largest deviations for $(u=1,s=0)$ are at the beginning and at the end of the sequence which affect the ON symbol edges. On the contrary, with $(u=N_{\rm seg}-1,s=0)$, the largest deviation is in the middle of the sequence, setting the highest envelope fluctuation right in the middle of the ON symbol. By selecting a different shift $s=N_{\rm seg}/2$ for $u=(N_{\rm seg}-1)$, the deviation of the phase increment then overlaps with case $(u=1,s=0)$, and thus both ZC sequences create almost the same OOK envelope shape. Finally, the corresponding powers of subcarrier coefficients for these signals are shown in Fig. 11(f). Every two coefficients are equal to zero which is due to $B[k]$ being equal here to the repetition of the first column in the second table of Table I. Every fourth coefficient has the same power as shown in Corr. 5, accounting for half of the total power as stated in Lem. 4. Other coefficients’ power fluctuates according to (28).
V-B Concentrating ON Symbol Energy
@@
…for improving robustness against timing error: It was observed in [2] that while OOK-modulated LP-WUS is robust against frequency synchronization error, it is sensitive to timing error. The larger $N_{\rm bit}$ is, the smaller the OOK symbols are, and the less maximum timing error can be tolerated. LRs are expected to use low-cost oscillator with low-precision frequency synchronization, resulting in timing clock drifting. Since LP-WUS is a sporadic message based on needs, its precise transmission time and thus resulting timing drift from previous synchronization acquisition cannot be known at LR. We will model this timing error as uniformly distributed in range $[-\tau_{\rm err},\,\tau_{\rm err}]$ where $\tau_{\rm err}$ is the maximum expected timing error around current receiver timing, given for example from the synchronization signal periodicity.
@@ V-B3 Concentrated Receiver Window
VI Numerical Evaluations
In this section, the benefits of shaped OOK designs discussed in Section V compared to rectangular OOK designs discussed in Section IV are shown by numerical evaluations in the context of 3GPP simulations assumptions for LP-WUS.
@@ VI Numerical Evaluations
Figure 13: BER performance comparison between different rectangular OOK designs and shaped-OOK designs with spreaded sprectrum.
Fig. 13 shows the bit errot rate (BER) for different OOK designs under the assumption of perfect synchronization. The worst-performing curve corresponds to a plain DFT-s-OFDM modulation where bits are only trivially spread using an all-one sequence as $r[m]=1$, without additional processing (i.e. $L=\Phi=0$, $N_{\rm sc}=N_{\rm pulse}$ and no FDSS as $W[k]=1$). This method performs poorly because it fails to provide a good energy split between ON and OFF symbols, leading to an envelope shape similar shape to the case of $\Phi=0$ in Fig.6(a). The rectangular-like OOK signals provide notable detection improvement, as they achieve an excellent energy split between ON and OFF symbols. Three different variations are considered: i) (18) without FDSS; ii) (18) with FDSS using $\beta=4$ and a time-shifting of half a pulse; and iii) LS waveform discussed in Sec. IV-C. These three rectangular-like OOK designs perform equivalently, showing that envelope fluctuation inside ON symbols has no effect on performance. However, the performance of these designs is limited by their concentrated spectrum. By applying frequency repetition on top of rectangular-like design based on (18) as discussed in Sec. V-A1, the performance improves significantly but saturates after 3 repetitions. Here, 2, 3 and 4 repetitions are achieved by setting $N_{\rm pulse}=68$, $44$ and $32$, respectively. The best performance is obtained using a ZC overlaid sequence as discussed in Sec. V-A3. In this case, $L=\Phi=0$, $N_{\rm sc}=N_{\rm pulse}$ and there are no FDSS. The length of the ZC sequence is $N_{\rm seg}=33$, with two roots, $u=1$ and $17$, considered. Both roots perform similarly, though the middle root $u=17$ performs slightly better. This can be attributed to a slightly better ON/OFF energy split with less OFF symbol fluctuation.
@@
Figure 14: BER performance comparison between normal OOK and concentrated OOK.
@@ VII Conclusions
This paper studied a general framework for DFT-s-OFDM based OOK modulation, motivated by the recent adoption of OOK wake-up signals in modern wireless standards. Equivalent frequency-domain implementations were identified. Rectangular-like OOK signals were designed, including with a zero-DC property and equivalent as with the LS precoding method. Finally, shaping methods useful for robustness against frequency-selective fading and timing error were analyzed and numerically evaluated. The results presented herein provide background and further insights into key signal design considerations of 3GPP LP-WUS study. 5G-Advanced LP-WUS OOK signal design, whose details are currently under specification by 3GPP, is expected to be reused for other future IoT technologies such as Ambient-IoT.
@@
$D[k]$ $=$ $\sum_{m=0}^{N_{\rm pulse}-1}b\left[\lfloor m/N_{\rm seg}\rfloor\right]r[m]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}km}$ (33)
$=$ $\sum_{m=0}^{N_{\rm bit}-1}\sum_{l=mN_{\rm seg}}^{(m+1)N_{\rm seg}-1}b\left[\lfloor l/N_{\rm seg}\rfloor\right]r[l]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}kl}$
$=$ $\sum_{m=0}^{N_{\rm bit}-1}b\left[m\right]\sum_{l=mN_{\rm seg}}^{(m+1)N_{\rm seg}-1}r[l]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}kl}.$
Now assuming $r[l]=e^{{\rm j}\Phi l}r_{0}[m-N_{\rm seg}l_{m}]$, we can apply a change a variable in the inner sum to get
$D[k]=\!\!\!\sum_{m=0}^{N_{\rm bit}-1}b\left[m\right]\!\!\!\sum_{l=0}^{Nseg-1}\!\!\!e^{{\rm j}\Phi(l+mN_{\rm seg})}r_{0}[l]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}k(l+mN_{\rm seg})}$
$=\!\!\!\sum_{m=0}^{N_{\rm bit}-1}\!\!\!b\left[m\right]e^{{\rm j}\Phi mN_{\rm seg}-{\rm j}\frac{2\pi}{N_{\rm pulse}}kmN_{\rm seg}}\!\!\!\sum_{l=0}^{N_{\rm seg}-1}\!\!\!e^{{\rm j}\Phi l}r_{0}[l]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}kl}$
$=\left(\sum_{m=0}^{N_{\rm bit}-1}\!\!\!b\left[m\right]e^{-{\rm j}\frac{2\pi m\left(k-\Phi N_{\rm pulse}\right)}{N_{\rm bit}}}\right)\!\!\!\sum_{l=0}^{N_{\rm seg}-1}\!\!\!r_{0}[l]e^{-{\rm j}\frac{2\pi l\left(k-\frac{\Phi N_{\rm pulse}}{2\pi}\right)}{N_{\rm pulse}}}$
$=R_{0}\left[k-\frac{\Phi N_{\rm pulse}}{2\pi}\right]B\left[k-\frac{\Phi N_{\rm pulse}}{2\pi}\right]$ (34)
where $R_{0}[f]$ and $B[f]$ are the interpolated DFT coefficients of the overlaid sequence and the bit sequence, respectively.
@@
$=$ $\sum_{n=0}^{N_{\rm bo}-1}e^{-{\rm j}2n\frac{\pi k}{N_{\rm bo}}}\left(\overline{b_{o}[n]}+b_{o}[n]e^{-{\rm j}\frac{\pi k}{N_{\rm bo}}}\right)$
@@
$\angle\frac{h\left[n-\frac{N_{\rm fft}}{N_{\rm pulse}}(m+1)\right]}{h\left[n-\frac{N_{\rm fft}}{N_{\rm pulse}}m\right]}=-\frac{(N_{\rm sc}-1)}{N_{\rm pulse}}\pi+\theta[n]$ (38)
@@
The function $\theta[n]=\{0\text{ or }\pi\}$ corresponds to a sign difference between the real part of the pulses, and changes as a function of $n$. In the case of $N_{\rm pulse}=N_{\rm sc}$ and no FDSS, one can verify from (11) $\theta[n]=0$ for the samples between the two neighboring main lobe peaks, and as $N_{\rm pulse}$ is decreased, $\theta[n]=0$ in a smaller range of samples around the crossing of the main lobes. Numerical investigation confirms it is also the case with common FDSS windows. So in general $\theta[n]=0$ where pulse combining should be the most avoided, which leads to

### v2

@@ V-A1 Frequency-Repetition
Starting from the rectangular-like OOK design discussed so far, a generic method to achieve higher frequency diversity is to repeat its subcarrier coefficients on a wider bandwidth. This can be achieved in (7) by selecting $N_{\rm pulse}\ll N_{\rm sc}$, i.e. modulating less, and more spread-apart, DFT-s-OFDM pulses, leading inevitably to a comb-like signal shape. In order to guarantee a good ON/OFF energy split, we first apply (20) to the intermediate narrow band of $N_{\rm pulse}$ subcarriers, i.e. setting $\Phi=\frac{\pi(2L+N_{\rm pulse}-1)}{N_{\rm pulse}}$, before repeating in (7) to populate the $N_{\rm sc}$ subcarriers. An example of the obtained waveforms and power distribution over subcarrier indices is shown on Fig. 10 with $N_{\rm sc}=144$ and $N_{\rm bit}=4$. No FDSS is used here other than a linear phase for time-shifting the signal by half a pulse. As it can be seen the ON symbols cannot be flattened since the main lobes of neighboring pulses simply do not overlap, but by selecting a proper $(\Phi,L)$, the OFF symbols can be guaranteed to be close to zero. Similarly to Fig. 6, variations over $(\Phi,L)$ can, in fact, be tolerated with little effect on the signal shape until high envelope lobes are created in the OFF symbols, similar to the green curve of Fig. 8. The drawback of this approach is as $N_{\rm pulse}\to N_{\rm bit}$, the signal becomes less robust to lower sampling rate if considered, and the PAPR of the overall OFDM transmission may increase even beyond what it is already with QAM symbols.
@@ V-A2 Phase Scrambling in Overlaid Sequence
Methods for varying the phase of the overlaid sequence $r_{0}[k]$ in order to flatten the spectrum were considered in [31]. While by scrambling the phase of the spreading sequence, the envelope shape of the ON symbol may become less controlled, the ON/OFF energy split can still be expected to remain good as it is unlikely that many distant sidelobes in the OFF symbols would coherently combine (cf. $O_{l}[n]$ in (13) for samples outside of $[lN_{\rm seg},\,(l+1)N_{\rm seg}-1]$).
As shown from Corr. 1, the FD power distribution depends on two sequences: $R_{0}[k]$ from the overlaid sequence, and $B[k]$ from the bits which inherently has a repetition structure such that its power is already well-distributed across the band. The FD power distribution of DFT-s-OFDM-based OOK satisfies then the following property shown in Appendix -E.
@@
With Manchester coding [^6] Without Manchester coding, $E[|B[k]|^{2}]=\frac{N_{\rm bit}(N_{\rm bit}+1)}{4}$ for $k=(0\,{\rm mod}\,N_{\rm bit})$ and $E[|B[k]|^{2}]=\frac{N_{\rm bit}}{4}$ otherwise; and more than half, precisely $\frac{(N_{\rm bit}+1)}{2N_{\rm bit}}$, of the average sum power is allocated on $\mathcal{C}$., half of the sum power of $\{D_{0}[k]\}_{k=0}^{N_{\rm pulse}-1}$ is always allocated on a regular comb of subcarrier indices $\mathcal{C}=\{k^{\prime}N_{\rm bit}\}_{k^{\prime}=0}^{N_{\rm seg}-1}$. On average, the power distribution is given by $E[|D_{0}[k]|^{2}]=|R_{0}[k]|^{2}E[|B[k]|^{2}]$ with
$E[|B[k]|^{2}]=\begin{cases}\frac{N_{\rm bit}^{2}}{4}\quad\quad{\rm for}\;\;(k\,{\rm mod}\,N_{\rm bit})=0\\ \frac{N_{\rm bit}}{4}\left(1-\cos\frac{2\pi k}{N_{\rm bit}}\right)\quad{\rm otherwise}\end{cases}.$ (29)
While having half of the power allocation on this regular comb of subcarriers $\mathcal{C}$ is independent of $R_{0}[k]$, the power level fluctuation on $\mathcal{C}$ depends on it. Meanwhile, the power level of $R_{0}[k]$ on $\mathcal{C}$ is directly given by the DFT of $r_{0}[l]$, $R_{0}^{\prime}[k]=\sum_{l=0}^{N_{\rm seg}-1}r_{0}[l]e^{-{\rm j}\frac{2\pi}{N_{\rm seg}}kl}$, since $R_{0}[k]$ is an $N_{\rm bit}$-upsampled and sinc-interpolated version of $R_{0}^{\prime}[k]$ given by
$\!\!\!R_{0}[k]=\begin{cases}R^{\prime}_{0}[k/N_{\rm bit}]\quad\quad\quad{\rm for}\;\;(k\,{\rm mod}\,N_{\rm bit})=0\\ \displaystyle\frac{1}{N_{\rm seg}}\!\!\!\sum_{h=0}^{N_{\rm seg}-1}\!\!\!R^{\prime}_{0}[h]I[k-N_{\rm bit}h]\quad{\rm otherwise,}\!\!\!\end{cases}$ (30)
with $I[k]=e^{-{\rm j}\frac{\pi(N_{\rm seg}-1)k}{N_{\rm pulse}}}\frac{\sin\left(\frac{\pi k}{N_{\rm bit}}\right)}{\sin\left(\frac{\pi k}{N_{\rm pulse}}\right)}$, see Appendix -F.
So by combining Lem. 4 with (30), the average power of the subcarrier coefficients $E[|D_{0}[k]|^{2}]$ is completely characterized; from which one sees that half of the average sum power can be directly controlled by the overlaid sequence $r_{0}[l]$ to be distributed with a constant level over the band.
@@
Randomizing the phase of $r_{0}[m]$ has been proposed in [31] to ensure a well-spread energy along the band. Indeed, this approach equalizes on average the power of its DFT coefficients since it leads to $E[|R_{0}[k]|^{2}]=N_{\rm seg},\,\forall k$ , which combined with Lem. 4 gives
@@
With random overlaid sequence such that $E[r_{0}[m]r_{0}^{*}[m^{\prime}]]=\delta_{m,m^{\prime}}$, half of the average sum power of $\{D_{0}[k]\}_{k=0}^{N_{\rm pulse}-1}$ is equally distributed on a $N_{\rm bit}$-spaced comb.
The drawback of this approach is that the signal is not controlled anymore, and therefore, its knowledge could not be exploited by receivers with higher-end capabilities.
@@
The approach retained for 5G LP-WUS is to use a ZC sequence [^7] For the purpose also of carrying the bits by overlaid sequence selection, in parallel to OOK, enabling improved performance for higher-end LRs. which is a constant-envelope zero autocorrelation (CAZAC) sequence. A well-known property of a CAZAC sequence is that its DFT is also of constant magnitude [32], such that $|R_{0}^{\prime}[h]|^{2}=N_{\rm seg},\,\forall h$ in (30). Combined with Lem. 4 this gives the same result as in Corr. 4 but for each OFDM symbol without averaging.
@@
With a CAZAC overlaid sequence, half of the sum power of $\{D_{0}[k]\}_{k=0}^{N_{\rm pulse}-1}$ is equally distributed on a regular $N_{\rm bit}$-spaced comb.
@@
Comparing to (20) for flattening ON symbols, the closer $\Delta_{\theta}[m]$ would be constant and close to $\Phi\approx\pi$ in this case, the better the DFT-s-OFDM pulses are expected to constructively combine. The deviations of this linear phase increment $\Delta_{\theta}$ from a constant increment of $\pi$ are shown on Fig. 11(e). With roots closer to $(0\mod N_{\rm seg})$, e.g. $u=1$ or $u=(N_{\rm seg}-1)$, the increment changes slower than with ‘middle roots’, e.g. $u=(N_{\rm seg}/2+1)$, which is better for pulse combining. The largest deviations for $(u=1,s=0)$ are at the beginning and at the end of the sequence which affect the ON symbol edges. On the contrary, with $(u=N_{\rm seg}-1,s=0)$, the largest deviation is in the middle of the sequence, setting the highest envelope fluctuation right in the middle of the ON symbol. By selecting a different shift $s=N_{\rm seg}/2$ for $u=(N_{\rm seg}-1)$, the deviation of the phase increment then overlaps with case $(u=1,s=0)$, and thus both ZC sequences create almost the same OOK envelope shape. Finally, the corresponding powers of subcarrier coefficients for these signals are shown in Fig. 11(f). Every two coefficients are equal to zero which is due to $B[k]$ being equal here to the repetition of the first column in the second table of Table I. Every fourth coefficient has the same power as shown in Corr. 5, accounting for half of the sum power as stated in Lem. 4. Other coefficients’ power fluctuates according to (30).
V-B TD Shaping for Improved Robustness Against Timing Error
It was observed in [2] that while OOK-modulated LP-WUS is robust against frequency synchronization error, it is sensitive to timing error. The larger $N_{\rm bit}$ is, the smaller the OOK symbols are, and the less maximum timing error can be tolerated. LRs are expected to use low-cost oscillator with low-precision frequency synchronization, resulting in timing clock drifting. Since LP-WUS is a sporadic message based on needs, its precise transmission time and thus resulting timing drift from previous synchronization acquisition cannot be known at LR. We will model this timing error as uniformly distributed in range $[-\tau_{\rm err},\,\tau_{\rm err}]$ where $\tau_{\rm err}$ is the maximum expected timing error around current receiver timing, given for example from the synchronization signal periodicity.
@@ V-B3 Concentrated Receiver Window
VI 5G LP-WUS Performance Evaluations
In this section, the benefits of shaped OOK designs discussed in Section V compared to rectangular OOK designs discussed in Section IV are evaluated in the context of 3GPP simulations assumptions for LP-WUS.
@@
Figure 13: BER performance comparison between different rectangular-like OOK designs and shaped-OOK designs with spread power sprectrum.
VI-A Performance Gain of OOK with FD Power Spreading over Rectangular-like OOK
Fig. 13 shows the bit error rate (BER) for different OOK designs under the assumption of perfect synchronization. The worst-performing curve corresponds to a plain DFT-s-OFDM modulation where bits are only trivially spread using an all-one sequence as $r[m]=1$, without additional processing (i.e. $L=\Phi=0$, $N_{\rm sc}=N_{\rm pulse}$ and no FDSS as $W[k]=1$). This method performs poorly because it fails to provide a good energy split between ON and OFF symbols, leading to an envelope shape similar to the case of $\Phi=0$ in Fig. 6(a). The rectangular-like OOK signals provide notable detection improvement, as they achieve an excellent energy split between ON and OFF symbols. Three different variations are considered: i) (20) without FDSS; ii) (20) with FDSS using $\beta=4$ and a time-shifting of half a pulse; and iii) LS waveform discussed in Sec. IV-C. These three rectangular-like OOK designs perform equivalently, showing that envelope fluctuation inside ON symbols has no effect on performance. However, the performance of these designs is limited by their concentrated spectrum. By applying frequency repetition on top of rectangular-like design based on (20) as discussed in Sec. V-A1, the performance improves significantly but saturates after 3 repetitions. Here, 2, 3 and 4 repetitions are achieved by setting $N_{\rm pulse}=68$, $44$ and $32$, respectively. The best performance is obtained using a ZC overlaid sequence as discussed in Sec. V-A3. In this case, $L=\Phi=0$, $N_{\rm sc}=N_{\rm pulse}$ and there are no FDSS. The length of the ZC sequence is $N_{\rm seg}=33$, with two roots, $u=1$ and $17$, considered. Both roots perform similarly, though the middle root $u=17$ performs slightly better. This can be attributed to a slightly better ON/OFF energy split with less OFF symbol fluctuation.
VI-B Performance Gain and Optimization of Concentrated OOK
(a) Comparison between normal and concentrated OOK.
(b) Different numbers of guard pulses.
(c) Concentration at Tx versus Rx.
Figure 14: BER evaluations of concentrated OOK with timing offset.
@@ VII Conclusions
This paper studied a general framework for DFT-s-OFDM-based OOK modulation, motivated by the recent adoption of OOK wake-up signals in modern wireless standards. Equivalent frequency-domain implementations were identified. Rectangular-like OOK signals with very flat OOK symbols and sharp edge transition were obtained, possibly with a zero-DC property, and recovering LS approach as a special case. Finally, frequency-domain and time-domain shaping methods useful for robustness against frequency-selective fading and timing error, respectively, were analyzed and evaluated. The results presented herein provide background and further insights into key signal design considerations of 3GPP LP-WUS study. 5G-Advanced LP-WUS OOK signal design, whose details have been recently specified by 3GPP, is expected to be reused for other future IoT technologies as already done for 5G-Advanced Ambient-IoT.
Acknowledgments
The first author is grateful to Branislav M. Popović for valuable discussions on LP-WUS and his reviews of preliminary internal reports. The authors also thanks their 3GPP standardization colleagues Xue Yifan, Zhou Han, Li Qiang, Long Yi and Matthew Webb for fruitful discussions related to LP-WUS and Ambient-IoT.
@@ Acknowledgments
In a general setting of a bit spreading by $N_{\rm seg}=N_{x}/N_{\rm bit}$ where $N_{x}$ is only restricted to be a factor of $N_{\rm bit}$ such that $N_{\rm seg}$ is an integer, the modulation symbols $d[m]$ are then defined as in (5) but for $m=0,\ldots,N_{x}-1$, and the pre-processed subcarrier coefficients are $D[k]=\sum_{m=0}^{N_{x}-1}d[m]e^{-j\frac{2\pi}{N_{x}}km}$; the following result holds.
Lemma 5.
$D[k]=\sum_{l=0}^{N_{\rm bit}-1}\tilde{R}_{l}[k]b[l]e^{-{\rm j}\frac{2\pi}{N_{\rm bit}}kl}$ (35)
where the $\tilde{R}_{l}[k]=\sum_{m=0}^{N_{\rm seg}-1}\tilde{r}_{l}[m]e^{-{\rm j}\frac{2\pi}{N_{x}}km}$ is the DFT of the $l$th overlaid sequence zero-padded to length $N_{x}$.
Proof.
$D[k]$ $=$ $\sum_{m=0}^{N_{x}-1}b\left[l_{m}\right]r[m]e^{-{\rm j}\frac{2\pi}{N_{x}}km}$ (36)
$=$ $\sum_{l=0}^{N_{\rm bit}-1}\sum_{m=lN_{\rm seg}}^{(l+1)N_{\rm seg}-1}b\left[l_{m}\right]r[m]e^{-{\rm j}\frac{2\pi}{N_{x}}km}$
$=$ $\sum_{l=0}^{N_{\rm bit}-1}b\left[l\right]\!\!\!\sum_{m=lN_{\rm seg}}^{(l+1)N_{\rm seg}-1}\!\!\!\tilde{r}_{l_{m}}[m\;{\rm mod}\;N_{\rm seg}]e^{-{\rm j}\frac{2\pi}{N_{x}}km}$
$=$ $\sum_{l=0}^{N_{\rm bit}-1}b\left[l\right]\sum_{m=0}^{N_{\rm seg}-1}\tilde{r}_{l}[m]e^{-{\rm j}\frac{2\pi}{N_{x}}k(m+lN_{\rm seg})}.$
which leads to (35) since $N_{\rm seg}=N_{x}/N_{\rm bit}$. ∎
An interpretation of the result above is that the sequence $\tilde{R}_{l}[k]$ creates an OOK symbol for the $l$th bit, and these OOK symbols are time-shifted and multiplexed together within the OFDM symbol in (35) according to the corresponding bit positions and values.
Under the assumption of (14): $\tilde{r}_{l}[m]=e^{{\rm j}\Phi(m+lN_{\rm seg})}r_{0}[m]$, DFTs of the overlaid sequences can be expressed as
$\tilde{R}_{l}[k]=e^{{\rm j}\Phi lN_{\rm seg}}R_{0}\left(k-\frac{\Phi N_{x}}{2\pi}\right)$ (37)
where $R_{0}(f)=\sum_{m=0}^{N_{\rm seg}-1}r_{0}[m]e^{-{\rm j}\frac{2\pi mf}{N_{x}}}$ is the interpolated DFT of the zero-padded common overlaid sequence $r_{0}[n]$, independent of the bit index $l$. Then, (35) can be rewritten as the product of two DFTs:
$D[k]=R_{0}\left(k-\frac{\Phi N_{x}}{2\pi}\right)B\left(k-\frac{\Phi N_{x}}{2\pi}\right)$ (38)
where $B(f)=\sum_{l=0}^{N_{\rm bit}-1}b\left[l\right]e^{-{\rm j}\frac{2\pi lf}{N_{\rm bit}}}$ is the interpolated DFT coefficients of the bit sequence.
Lem. 1 follows directly from (38) with $N_{x}=N_{\rm pulse}$.
Cor. 2 follows from (38) by taking $N_{x}=N_{\rm pulse}$ and the assumption of a common all-one overlaid sequence, i.e. $r_{0}[m]=1,\,\forall m$ as then by using the exponential sum formula one gets $R_{0}(f)=e^{-{\rm j}\pi f\frac{(N_{\rm seg}-1)}{N_{x}}}\frac{\sin\left(\frac{\pi}{N_{\rm bit}}f\right)}{\sin\left(\frac{\pi}{N_{x}}f\right)}$ for $f\neq 0\ {\rm mod}\ N_{x}$, and $R_{0}(f)=N_{\rm seg}=N_{x}/N_{\rm bit}$ otherwise.
Lem. 3 follows similarly by identifying (25) with Lem. 5 and then (38) where $\Phi=0$, $N_{x}=N_{\rm fft}$, $r_{0}[m]=1,\,\forall m$, and simplification of $R_{0}(f)$ by exponential sum formula. The final result follows by applying the shift and truncation of (24).
@@
$=$ $\sum_{n=0}^{N_{\rm bo}-1}e^{-{\rm j}2n\frac{\pi k}{N_{\rm bo}}}\left(\overline{b_{o}[n]}+b_{o}[n]e^{-{\rm j}\frac{\pi k}{N_{\rm bo}}}\right).$
@@
$\angle\frac{h\left[n-\frac{N_{\rm fft}}{N_{\rm pulse}}(m+1)\right]}{h\left[n-\frac{N_{\rm fft}}{N_{\rm pulse}}m\right]}=-\frac{(N_{\rm sc}-1)}{N_{\rm pulse}}\pi+\theta[n].$ (42)
@@
The function $\theta[n]=\{0\text{ or }\pi\}$ corresponds to a sign difference between the real part of the pulses, and changes as a function of $n$. In the case of $N_{\rm pulse}=N_{\rm sc}$ and no FDSS, one can verify from (11) that $\theta[n]=0$ for the samples between the two neighboring main lobe peaks, and as $N_{\rm pulse}$ is decreased, $\theta[n]=0$ in a smaller range of samples around the crossing of the main lobes. Numerical investigation confirms it is also the case with common FDSS windows. So in general $\theta[n]=0$ where pulse combining should be the most avoided, which leads to